OpenStudy (anonymous):

if the perimeter of a rectangle is 68cm. if the diagonal is 6cm find the dimentions of the rectangle. solved using simultanious equations ?

OpenStudy (anonymous):

perimeter of a rectangle is given by, 2L +2B= 68 , where L is length and B is breadth. diagonal can be formed by using Pythagoras theorem, 6^2 = L^2 + B^2 36 = L^2 + B^2 From 2L +2B= 68, Rearrange to form, L = (68-2B)/2 = 34 -B Substitute L = 34-B to the diagonal equation 36 = (34 -B)^2 + B^2 36 = 1156- 68B + B^2 + B^2 2B^2 - 68B + 1120 = 0 B^2 - 34B + 560 =0 And here im stuck, because there is no solution for B I believe there is an error in the question. Pretty sure the diagonal cannot be just 6cm

OpenStudy (anonymous):

sorrry the diagonal is 26 mmy keyboard didnt work properly, sorry for inconvenience

OpenStudy (anonymous):

Damn. Continuing from the substitution. 26^2 = (34 -B)^2 + B^2 676 = 1156- 68B + B^2 + B^2 2B^2 - 68B + 480 = 0 B^2 - 34B + 240 = 0 Solving quadratic equation, (B-10)(B-24) = 0 B = 10 or B= 24 Remember that 2L +2B= 68 L +B = 34 When B = 10 L +10 = 34 L = 24 When B = 24 L + 24 = 34 L = 10 This is rejected because length must be longer than breadth. Thus length of rectangle is 24 while breadth is 10.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np. hope you understand the workings.

OpenStudy (anonymous):

yeah thanks theyre nice and clear